OpenStudy (anonymous):

can someone let me know if I understand this correctly? 17/sqrt(17) multiply by sqrt(17)/sqrt(17) in denominator= sqrt(17) times sqrt(17) equals 17. The 17 in the numerator and the 17 in the denominator cancel each other out to equal sqrt(17)/1 or just sqrt(17).

mathslover (mathslover):

I can not understand the second para correctly but YES the ifrst step is right

mathslover (mathslover):

*first \[\large{\frac{17}{\sqrt{17}}=\frac{17*\sqrt{17}}{\sqrt{!7}*\sqrt{17}}}\]

OpenStudy (saifoo.khan):

Your understood it correctly! @a.galvan2

OpenStudy (jiteshmeghwal9):